OpenStudy (anonymous):

I have not done asymtotes since January and I threw away my notes when school ended :/ x-3/x+2 how do you figure out the vertical and horizontal asymptote

OpenStudy (agent0smith):

Vertical: Look for values of x that'll make the denominator equal zero. The denominator is x+2... find x that'll make it zero, that's your vertical asymptote, x = ... Horizontal: Look at the highest powers of x in the numerator and denominator... they are x and x. What is x/x equal to? That's your horizontal asymptote, y = ...

OpenStudy (anonymous):

I remember a little of the horizontal, do I just see which one higher than the other?

OpenStudy (agent0smith):

Yes, basically. y = x/x = ...?

OpenStudy (anonymous):

so x/x is equal to 1

OpenStudy (agent0smith):

Correct. H.A. is y=1.

OpenStudy (anonymous):

when will the horizontal = 0 or none

OpenStudy (agent0smith):

H.A. will be y = 0 when the highest power of x on the top is smaller than the bottom. Eg \[\Large y = \frac{ x }{ x^2 }\] H.A. won;'t exist when it's the other way around... highest power on the top is bigger than the bottom eg. \[\Large y = \frac{ x^2 }{ x }\] Well it will have a "slant" asymptote, but you may not have to worry about those yet.

OpenStudy (anonymous):

I have this problem where its (x+6)(x-6)/x^2+9 would the H.A. 1 or none

OpenStudy (agent0smith):

well the top part has an x^2 term after you expand out (x+6)(x-6) to find the H.A. \[\Large y = \frac{ x^2 }{x^2 } = ...?\]